OpenStudy (anonymous):

While a person is walking, his arms swing through approximately a 45.0 angle in 0.520 . As a reasonable approximation, we can assume that the arm moves with constant speed during each swing. A typical arm is 70.0 long, measured from the shoulder joint. What is the acceleration of a 0.700 gram drop of blood in the fingertips at the bottom of the swing?

OpenStudy (anonymous):

I got a 11.396 \[m/s^{2}\] as an answer but it doesn't seem to be right? I would really appreciate it if someone is willing to help~

OpenStudy (anonymous):

I did ω=θ/t and got 1.51 rad/sec then did a = g + Rω2 = 11.396

OpenStudy (anonymous):

What's the 0.520? Seconds?

OpenStudy (anonymous):

Yes it is seconds

OpenStudy (anonymous):

Is the middle of his swing perpendicular to the ground?

OpenStudy (anonymous):

I would presume so, yes, because the question has not stated otherwise~

OpenStudy (anonymous):

Alright, give me some time to do it out. Do you just want me to corroborate your answer or give you hints? Because I'd prefer to not actually calculate. XD

OpenStudy (anonymous):

Hints are fine and more preferable, I would like to work out the question without getting the answer so I can try to figure it out myself as well. Thank you~

OpenStudy (anonymous):

Velocity is constant (thank God) and the velocity vector at the bottom of the swing is directly parallel to the ground. Assuming no drag, the only F=ma acting on the droplet of blood is vertical F=mg.

OpenStudy (anonymous):

Unless I'm doing something wrong, no calculations are necessary. :o I sense a trick question.

OpenStudy (anonymous):

Hmm...let me see what I get if your right

OpenStudy (anonymous):

Well, it should be acceleration of gravity...

OpenStudy (anonymous):

Aw...got it wrong apparently the answer seems to be 1.60\[m/s^{2}\]

OpenStudy (anonymous):

Don't know where that came from...

OpenStudy (anonymous):

Huh, that's interesting. I must've made a mistake or misread the question. My bad. But I've no idea what I'm doing wrong. >.> Better wait for someone else to step in.

OpenStudy (anonymous):

That's what happened to me as well, I thought I was right but...oh well

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Pendulum Wikipedia agrees with me. ಠ_ಠ

OpenStudy (anonymous):

Are you sure you typed out the full question?

OpenStudy (anonymous):

Yes, the only thing I left out was the other 2 questions which I think I can get myself

OpenStudy (anonymous):

http://upload.wikimedia.org/wikipedia/commons/2/24/Oscillating_pendulum.gif This makes it pretty clear that the velocity vector is parallel to the ground, and the acceleration vector is perpendicular--however, we're told that the arm swinging is at a constant velocity, so there's no acceleration.

OpenStudy (anonymous):

I agree, but if that would hold true, either the answer is wrong or we're just missing something

OpenStudy (jamesj):

So there are three forces acting: gravity, the normal force against the blood of the arm and the centripetal force. As the motion of the drop of blood shows, i.e., circular, the gravitational force plays no role; or rather, it is cancelled by the force of the blood vessel on it, the 'normal force'. The acceleration due to the centripetal force is \[ \omega^2 r = 1.51^2(0.7) = 1.6 \ m/s^2 \] which is your answer. This is a bit of a trick question, and to be honest, the last time a version of this question was asked, I missed the fact that the motion of the blood drop itself demonstrates that the net force is just the centripetal force.

OpenStudy (anonymous):

This is true. I forget that parabolic motion where the object is never in free fall can be modeled as a part of a circle, hence centripedal force. XD

OpenStudy (jamesj):

(You're making this slightly harder than it needs to be. There's no parabolic motion here. The blood drop is moving in a circle.)

OpenStudy (anonymous):

I'm sorry, pendulum motion. >.> Not parabolic.

OpenStudy (jamesj):

ah, I see.

OpenStudy (anonymous):

They both start with p, don't judge me XD